# Collisional Baryon-Dominated Dwarf Galaxies: A New Probe of Bursty Feedback and Dark Matter Physics

### v1

@@ Abstract
Discussion and Conclusion. Our controlled simulations show that energy injection weakens the progenitors’ baryonic binding, elevating BDDG formation through collisions. Repeated potential fluctuations, as expected from bursty stellar feedback, may reduce $|E_{\rm bind}|$ by only 15%, yet in two thirds of the simulations, the BDDG masses are enhanced by over 100%. In contrast, SIDM creates cores without lowering $|E_{\rm bind}|$ and has a negligible effect on BDDG masses.
@@
Observations of such systems therefore provide a powerful diagnostic of the energy injection mechanisms. Wide-field surveys (LSST, WFST, CSST, Roman, Euclid) will greatly expand catalogs of low-surface-brightness galaxies, identifying BDDG candidates through diffuse morphologies, old stellar populations, and bright globular clusters. However, optical imaging alone cannot confirm their baryon-dominated nature. Crucially, 21-cm surveys (FAST, MeerKAT, VLA) will provide gas fractions and kinematics, establishing whether baryons alone account for the observed dynamical support. Together, these facilities will enable systematic searches, constrain the frequency of BDDG formation, and map their mass spectrum.
Our study focuses on idealized collisions of field galaxies, and analogous processes could also occur in the collisional formation of TDGs, warranting further investigation. Extending our analysis to cosmological simulations will enable quantification of the BDDG population, determination of the fraction of associated systems, and exploration of the dependence on environment. Aside from energy injection by periodic potential fluctuations, the collisional formation of BDDGs is broadly sensitive to scenarios with weakened gravitational binding. In Warm Dark Matter models, particles possess intrinsically higher initial velocities, leading to shallower inner halo potentials and weaker gravitational binding [65]. Recently, Ref. [49] proposes a model featuring baryon feedback modulated by interactions between baryons and dark matter. Taken together, our results establish BDDG formation as a promising avenue for probing the interplay between feedback-regulated galaxy formation and the fundamental nature of dark matter.
[…]
@@
Acknowledgment This work is supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), the New Cornerstone Science Foundation through the XPLORER PRIZE, the China Manned Space Program (No. CMS-CSST-2025-A03), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092), the Postdoctoral Fellowship Program of CPSF (No. GZB20250738), and the Jiangsu Funding Program for Excellent Postdoctoral Talent (No. 2025ZB209).
@@ Appendix A Star formation in gas remnants
Figure 5: Comparison of star formation efficiency in cored and cuspy benchmarks. Maps of the $b_{T}$, which traces the star formation efficiency, are shown for the cuspy ($\gamma=1$, 2a, left) and cored ($\gamma=0.1$, 2b, right) benchmarks at $t=0.4$ Gyr (top) and $1.0$ Gyr (bottom). White circles mark the sites of the forming BDDGs. The cored ($\gamma=0.1$) halo develops a significantly brighter region, indicating higher $b_{T}$ values and enhanced star formation efficiency at the BDDG site. The cuspy ($\gamma=1$) halo, while forms four BDDGs, has substantially lower $b_{T}$ values.
Figure 6: Post-collision star formation. Mass evolution of the most massive BDDGs in the 2a ($\gamma=1$) and 2b ($\gamma=0.1$) simulations. In the weakened binding case ($\gamma=0.1$), efficient star formation yields a more massive BDDG after the collision, with continued activity beyond 2 Gyr steadily depleting the gas reservoir. For $\gamma=1$, star formation is much less efficient, leaving a substantial gas fraction even at 5 Gyr. In both scenarios, the total BDDG mass increases only modestly, from $11\times 10^{8}~\rm M_{\odot}$ ($2.3\times 10^{8}~\rm M_{\odot}$) to $15\times 10^{8}~\rm M_{\odot}$ ($2.6\times 10^{8}~\rm M_{\odot}$) in the $\gamma=0.1$ ($\gamma=1$) case.
Our simulations reveal a systematic trend in which star formation is more efficient when the gravitational binding is weakened. To demonstrate this, we compare the quantity $b_{T}$ for the $\gamma=1$ (left) and $\gamma=0.1$ (right) cases in Fig. 5 This parameter, motivated by the ratio of dynamical to free-fall times, increases with star formation efficiency. Following Ref. [66], we define
$b_{T}\equiv\frac{\Sigma_{\rm gas}}{\sigma_{v,\rm gas}^{2}+c_{s}^{2}}\propto\left(\frac{\tau_{\rm dyn}}{\tau_{\rm ff}}\right)^{2},$ (2)
@@
Figure 5 shows $b_{T}$ for the 2a benchmark ($\gamma=1$, left panels) and the 2b benchmark ($\gamma=0.1$, right panels) at $t=0.4$ Gyr (top) and $1.0$ Gyr (bottom). The eventual locations of BDDG formation are enclosed by white circles. In the central collision region, the $\gamma=0.1$ case exhibits systematically higher $b_{T}$, corresponding to enhanced star formation. Moreover, in the $\gamma=1$ run the gas remnants fragment into three BDDGs, whereas in the $\gamma=0.1$ run only a single remnant forms. Similar behavior is observed across multiple paired simulations.
Figure 6 compares the mass evolution of the most massive BDDGs in the 2a ($\gamma=1$) and 2b ($\gamma=0.1$) simulations. In the weakened gravitational binding scenario ($\gamma=0.1$), efficient star formation produces a larger BDDG mass after the collision. Continued star formation beyond 2 Gyr steadily depletes the gas, leading to a decreasing gas fraction. In the $\gamma=1$ case, star formation also reduces the gas fraction, but with much lower efficiency, leaving a substantial gas reservoir even at 5 Gyr. In both cases, the total BDDG mass changes only modestly, increasing from $11\times 10^{8}~\rm M_{\odot}$ ($2.3\times 10^{8}~\rm M_{\odot}$) to $15\times 10^{8}~\rm M_{\odot}$ ($2.6\times 10^{8}~\rm M_{\odot}$) in the $\gamma=0.1$ ($\gamma=1$) simulation.
@@ Appendix B Stability of SIDM cores
First, radiation pressure in the gaseous disk is subdominant compared to gravitational support; therefore, running the initial SIDM simulation without hydrodynamics does not break the equilibrium of the gas disk, allowing it to be transferred consistently into a subsequent hydrodynamical run. Second, the constant SIDM cross section per mass adopted here should be regarded as an effective constant cross section [67]. In velocity-dependent SIDM models consistent with observational constraints, the effect of SIDM during the collision can be neglected because the collision velocity is much larger than the typical particle velocities in the progenitor halos. Lastly, we assume that once SIDM cores form, the resulting cored profiles remain stable over relatively long timescales, even if SIDM interactions are subsequently turned off. This implies that SIDM halos reach a quasi-hydrostatic equilibrium, a point we explicitly verify with a toy simulation.
@@
Appendix C Modeling energy injection into dark matter halos
We model the impact of energy injection into dark matter halos using an analytic framework inspired by Ref. [28]. We initialize a spherical, isotropic NFW halo with mass $1.5\times 10^{10}~\rm M_{\odot}$ and concentration $c=14$, corresponding to the 1a and 1b benchmarks in Table 1. For each dark matter particle, we compute the total energy $E_{0}$ and specific angular momentum $j$. A sequence of impulsive potential fluctuations is modeled by rescaling the potential by a small fraction, $\Delta V/V_{0}=0.14$, at each iteration. The energy of each particle is updated using the phase-averaged kick at fixed $j$
$\Delta E=\frac{2nE_{0}}{(2+n)^{2}}\,\bigg(\frac{\Delta V}{V_{0}}\bigg)^{2},$ (3)
$p(r;E,j)$ $\propto$ $\frac{1}{\sqrt{E-V_{\rm eff}(r,j)}},$
$V_{\rm eff}(r,j)$ $=$ $V(r)+\frac{j^{2}}{2r^{2}},$
over all dark matter particles.
Appendix D Tabulated full simulation results
@@
We adopt the gas cooling and star formation module provided in the Gadget-4 code. Radiative cooling is implemented through tabulated rates for primordial gas, while stars form stochastically from dense gas following the multiphase ISM prescription of Springel & Hernquist (2003) [64]. Although our study does not focus on the detailed dynamical processes within UDGs, the efficiency of gas cooling plays a key role in regulating their stability: a lower cooling rate is generally expected to promote stability. To demonstrate that our conclusions are robust against variations in the cooling rate, we present the full simulation results in Table 2 for the default cooling rate and in Table 3 for a case with the cooling rate reduced by half. In these tables, we identify BDDGs as systems with $M_{*}(r<10{\rm kpc})>10^{6}M_{\odot}$, reporting their abundance as well as the mass content of the most massive BDDGs in each simulation.
@@
Figure 8 summarizes the reduced cooling case, in the same format as Fig. 4 in the main text. All key features, most notably the enhancement in BDDG mass and star formation efficiency, remain unaffected. A comparison of the numerical results across the two tables shows that reduced cooling produces more BDDGs with lower characteristic masses. This indicates a potential degeneracy between the effects of gas cooling and weakened gravitational binding. The degeneracy can be resolved by the observed abundance of UDGs, which provides a means to constrain gas cooling [68, 69, 70, 57].
@@
Both Fig. 4 and 8 present results for the most massive BDDGs. To further illustrate the robustness of our findings, Fig. 9 shows the total BDDG masses for the default cooling scenario. In all cases, the trend of enhanced BDDG masses in the weakened gravitational binding scenario remains unchanged.
@@
Figure 8: Characteristics of the most massive BDDGs in 14 paired simulations with reduced gas cooling. The bar chart summarizes results from contrasting simulations with inner slopes $\gamma=0.1$ and $\gamma=1$. Bars are arranged in ascending order of the BDDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $(E_{{\rm bind},\gamma=1}-E_{{\rm bind},\gamma=0.1})/E_{{\rm bind},\gamma=1}$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In contrast, the increase in the BDDG masses is predominantly positive. In 8 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$.
@@
Figure 9: Characteristics of the summed BDDG content for 15 paired simulations. The bar chart summarizes results from contrasting simulations with inner slopes $\gamma=0.1$ and $\gamma=1$. Bars are arranged in ascending order of the summed BDDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $(E_{{\rm bind},\gamma=1}-E_{{\rm bind},\gamma=0.1})/E_{{\rm bind},\gamma=1}$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In contrast, the increase in the BDDG masses is predominantly positive. In 4 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$.
Table 2: Galaxy collision simulations with Gadget-4: initial collision configurations and properties of the resulting dark-matter-dominated galaxies (DMDGs). The properties listed are: (1) dark matter halo mass $M_{200}$, (2) total gas mass, (3) initial relative collision velocity $v_{r}$, (4) gaseous disk scale radius, (5) concentration $c$, (6) $\gamma$, (7) the absolute value of the baryonic binding energy $|E_{\rm bind}|$, presented in unit $E_{0}\equiv(10^{12}\,M_{\odot}\,\rm km^{2}\,s^{-2})$, (8) number of resulting DMDGs with $M_{b}>10^{6}M_{\odot}$, (9) stellar mass of the most massive DMDG, and (10) its gas mass. All properties are measured within $10\,{\rm kpc}$ at $t=2.0\,{\rm Gyr}$.
BM $M_{200,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $|E_{\rm bind}|$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) $E_{0}$ ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)
@@
BM $M_{200,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $|E_{\rm bind}|$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) $E_{0}S$ ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)

### v2

@@ 5 Discussion and Conclusion
Our controlled simulations show that energy injection weakens the progenitors’ baryonic binding, elevating DMDG formation through collisions. Repeated potential fluctuations, as expected from bursty stellar feedback, may reduce $|E_{\rm bind}|$ by only 15%, yet in two thirds of the simulations, the DMDG masses are enhanced by over 100%. In contrast, SIDM creates cores without lowering $|E_{\rm bind}|$ and has a negligible effect on DMDG masses.
In the regime of high mass ratios and low collision velocities, close encounters between satellite and host galaxies have been proposed as a primary channel for producing DMDGs within host systems (J. Moreno et al., 2022; D. Yang et al., 2020; Y. Jing et al., 2019; G. Ogiya, 2018). In such interactions, however, tidal stripping acts on both the dark matter and baryonic components of the satellite. If baryonic feedback is too strong, it can make the baryons too diffuse and easily stripped, hindering the formation of DMDGs. In contrast, SIDM can generate cores without significantly expanding the stellar distribution, offering a more efficient mechanism for DMDG formation (D. Yang et al., 2020). The relative abundance of DMDGs in the field versus those in host halos thus provides a potential observational handle to distinguish between core formation driven by baryonic feedback and that induced by SIDM.
The distinct formation pathways of DMDGs lead to contrasting observational signatures. Systems produced through tidal stripping resemble ordinary satellites but typically exhibit lower surface brightness. In contrast, those formed from post-collision tidal debris contain a substantial stellar component with nearly uniform ages and metallicities, created in a short burst at the time of the encounter. Another class of baryon-dominated systems is the tidal dwarf galaxy, which forms from tidal-tail gas ejected during a strong interaction involving a massive disk (F. Bournaud & P. A. Duc, 2006; G. Gentile et al., 2007; S. Kaviraj et al., 2012; J. Zaragoza-Cardiel et al., 2024). These galaxies are typically embedded within tidal tails and generally lack bright globular clusters (J. Lee et al., 2024).
@@
Future observations of such systems will be highly informative. Wide-field surveys (LSST, WFST, CSST, Roman, Euclid) will significantly expand the catalogs of low-surface-brightness galaxies and enable the identification of DMDG candidates through their morphologies, stellar populations, and associated globular clusters. Crucially, 21-cm surveys (FAST, MeerKAT, VLA) will measure gas fractions and kinematics, confirming whether baryons alone can account for the observed dynamics.
Extending this analysis to more realistic and cosmological settings will be essential for quantifying the DMDG population both in the field and within host environments. However, meeting the simultaneous demands of high resolution and large cosmological volume remains challenging. Current high-resolution simulations that resolve individual supernova explosions are restricted to controlled simulations, and cosmological simulations treat bursty feedback as subgrid physics (E. Zhang et al., 2025; Y. Deng et al., 2024; J. Moreno et al., 2022; D. Nelson et al., 2019). Continued advances in simulations are therefore needed to clarify the formation processes and abundance of DMDGs.
@@
We thank Fangzhou Jiang and Hui Li for useful discussions. This work is supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), the National Science Foundation of China (No. 12588101), the New Cornerstone Science Foundation through the XPLORER PRIZE, the China Manned Space Program (No. CMS-CSST-2025-A03), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092), the Postdoctoral Fellowship Program of CPSF (No. GZB20250738), and the Jiangsu Funding Program for Excellent Postdoctoral Talent (No. 2025ZB209).
@@ Appendix A Star formation in gas remnants
Figure 5: Comparison of star formation efficiency in cored and cuspy benchmarks. Maps of the $b_{T}$, which traces the star formation efficiency, are shown for the cuspy ($\gamma=1$, 2a, left) and cored ($\gamma=0.1$, 2b, right) benchmarks at $t=0.4$ Gyr (top) and $1.0$ Gyr (bottom). White circles mark the sites of the forming DMDGs. The cored ($\gamma=0.1$) halo develops a significantly brighter region, indicating higher $b_{T}$ values and enhanced star formation efficiency at the DMDG site. The cuspy ($\gamma=1$) halo, while forms four DMDGs, has substantially lower $b_{T}$ values.
Our simulations reveal a systematic trend in which star formation is more efficient when the gravitational binding is weakened. To demonstrate this, we compare the quantity $b_{T}$ for the $\gamma=1$ (left) and $\gamma=0.1$ (right) cases in Figure 5 This parameter, motivated by the ratio of dynamical time to free-fall time, increases with star formation efficiency. Following A. K. Leroy et al. (2016), we define
$b_{T}\equiv\frac{\Sigma_{\rm gas}}{\sigma_{v,\rm gas}^{2}+c_{s}^{2}}\propto\left(\frac{\tau_{\rm dyn}}{\tau_{\rm ff}}\right)^{2},$ (A1)
@@
Figure 6: Post-collision star formation. Mass evolution of the most massive DMDGs in the 2a ($\gamma=1$) and 2b ($\gamma=0.1$) simulations. In the weakened binding case ($\gamma=0.1$), efficient star formation yields a more massive DMDG after the collision, with continued activity beyond 2 Gyr steadily depleting the gas reservoir. For $\gamma=1$, star formation is much less efficient, leaving a substantial gas fraction even at 5 Gyr. In both scenarios, the total DMDG mass increases only modestly, from $11\times 10^{8}~\rm M_{\odot}$ ($2.3\times 10^{8}~\rm M_{\odot}$) to $15\times 10^{8}~\rm M_{\odot}$ ($2.6\times 10^{8}~\rm M_{\odot}$) in the $\gamma=0.1$ ($\gamma=1$) case.
Figure 5 shows $b_{T}$ for the 2a benchmark ($\gamma=1$, left panels) and the 2b benchmark ($\gamma=0.1$, right panels) at $t=0.4$ Gyr (top) and $1.0$ Gyr (bottom). The eventual locations of DMDG formation are enclosed by white circles. In the central collision region, the $\gamma=0.1$ case exhibits systematically higher $b_{T}$, corresponding to enhanced star formation. Moreover, in the $\gamma=1$ run the gas remnants fragment into three DMDGs, whereas in the $\gamma=0.1$ run only a single remnant forms. Similar behavior is observed across multiple paired simulations.
Figure 6 compares the mass evolution of the most massive DMDGs in the 2a ($\gamma=1$) and 2b ($\gamma=0.1$) simulations. In the weakened gravitational binding scenario ($\gamma=0.1$), efficient star formation produces a larger DMDG mass after the collision. Continued star formation beyond 2 Gyr steadily depletes the gas, leading to a decreasing gas fraction. In the $\gamma=1$ case, star formation also reduces the gas fraction, but with much lower efficiency, leaving a substantial gas reservoir even at 5 Gyr. In both cases, the total DMDG mass changes only modestly, increasing from $11\times 10^{8}~\rm M_{\odot}$ ($2.3\times 10^{8}~\rm M_{\odot}$) to $15\times 10^{8}~\rm M_{\odot}$ ($2.6\times 10^{8}~\rm M_{\odot}$) in the $\gamma=0.1$ ($\gamma=1$) simulation.
@@ Appendix B Stability of SIDM cores
First, radiation pressure in the gaseous disk is subdominant compared to gravitational support; therefore, running the initial SIDM simulation without hydrodynamics does not break the equilibrium of the gas disk, allowing it to be transferred consistently into a subsequent hydrodynamical run. Second, the constant SIDM cross section per mass adopted here should be regarded as an effective constant cross section (D. Yang & H.-B. Yu, 2022). In velocity-dependent SIDM models consistent with observational constraints, the effect of SIDM during the collision can be neglected because the collision velocity is much larger than the typical particle velocities in the progenitor halos. Lastly, we assume that once SIDM cores form, the resulting cored profiles remain stable over relatively long timescales, even if SIDM interactions are subsequently turned off. This implies that SIDM halos reach a quasi-hydrostatic equilibrium, a point we explicitly verify with a toy simulation.
@@
Appendix C Tabulated full simulation results
@@
We adopt the gas cooling and star formation module provided in the Gadget-4 code. Radiative cooling is implemented through tabulated rates for primordial gas, while stars form stochastically from dense gas following the multiphase ISM prescription of V. Springel & L. Hernquist (2003). Although our study does not focus on the detailed dynamical processes within UDGs, the efficiency of gas cooling plays a key role in regulating their stability: a lower cooling rate is generally expected to promote stability. To demonstrate that our conclusions are robust against variations in the cooling rate, we present the full simulation results in Table 2 for the default cooling rate and in Table 3 for a case with the cooling rate reduced by half. In these tables, we identify DMDGs as systems with $M_{*}(r<10{\rm kpc})>10^{6}M_{\odot}$, reporting their abundance as well as the mass content of the most massive DMDGs in each simulation.
@@
Figure 8 summarizes the reduced cooling case, in the same format as Figure 4 in the main text. All key features, most notably the enhancement in DMDG mass and star formation efficiency, remain unaffected. A comparison of the numerical results across the two tables shows that reduced cooling produces more DMDGs with lower characteristic masses. This indicates a potential degeneracy between the effects of gas cooling and weakened gravitational binding. The degeneracy can be resolved by the observed abundance of UDGs, which provides a means to constrain gas cooling (Y. Rong et al., 2017; R. F. J. van der Burg et al., 2017; A. Di Cintio et al., 2017; M. G. Jones et al., 2018).
@@
Both Figure 4 and Figure 8 present results for the most massive DMDGs. To further illustrate the robustness of our findings, Figure 9 shows the total DMDG masses for the default cooling scenario. In all cases, the trend of enhanced DMDG masses in the weakened gravitational binding scenario remains unchanged.
@@
Figure 8: Characteristics of the most massive DMDGs in 14 paired simulations with reduced gas cooling. The bar chart summarizes results from contrasting simulations with inner slopes $\gamma=0.1$ and $\gamma=1$. Bars are arranged in ascending order of the DMDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $(E_{{\rm bind},\gamma=1}-E_{{\rm bind},\gamma=0.1})/E_{{\rm bind},\gamma=1}$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In contrast, the increase in the DMDG masses is predominantly positive. In 8 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$.
@@
Figure 9: Characteristics of the summed DMDG content for 15 paired simulations. The bar chart summarizes results from contrasting simulations with inner slopes $\gamma=0.1$ and $\gamma=1$. Bars are arranged in ascending order of the summed DMDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $(E_{{\rm bind},\gamma=1}-E_{{\rm bind},\gamma=0.1})/E_{{\rm bind},\gamma=1}$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In contrast, the increase in the DMDG masses is predominantly positive. In 4 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$.
Table 2: Galaxy collision simulations with Gadget-4: initial collision configurations and properties of the resulting dark-matter-dominated galaxies (DMDGs).
BM $M_{200,\,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $W/E_{\rm bind}$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)
@@
Note. — The properties listed are: (1) dark matter halo mass $M_{200}$, (2) total gas mass, (3) initial relative collision velocity $v_{r}$, (4) gaseous disk scale radius, (5) concentration $c$, (6) $\gamma$, (7) the absolute value of the baryonic binding energy $|E_{\rm bind}|$, presented in unit $E_{0}\equiv(10^{12}\,M_{\odot}\,\rm km^{2}\,s^{-2})$, (8) number of resulting DMDGs with $M_{b}>10^{6}M_{\odot}$, (9) stellar mass of the most massive DMDG, and (10) its gas mass. All properties are measured within $10\,{\rm kpc}$ at $t=2.0\,{\rm Gyr}$.
@@
BM $M_{200,\,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $W/E_{\rm bind}$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)